OpenStudy (anonymous):

2log(x^6y^5)+3log(1/x)-3log(y)

OpenStudy (zepp):

\[\large 2\log(x^6y^5)+3\log(\frac{1}{x})-3\log(y)\]

OpenStudy (lgbasallote):

condense?

OpenStudy (anonymous):

yes @zepp and @lgbasallote